Part A Determine the value of the equilibrium constant, Kgoal, for the reaction C(s)+1/2O2(g)+H2(g)<-->1/2CH3OH(g)+1/2CO(g), Kgoal=? by...

Part A
Determine the value of the equilibrium constant, Kgoal, for the reaction

C(s)+1/2O2(g)+H2(g)<-->1/2CH3OH(g)+1/2CO(g), Kgoal=?
by making use of the following information:

1. CO2(g)+3H2(g)<-->CH3OH(g)+H2O(g) K1 = 1.40 x 10^2

2. CO(g)+H2O(g)<-->CO2(g)+H2(g) K2 = 1.00 x10^5

3. 2C(s)+O2(g)<-->2CO(g) K3 = 2.10 x10^47
Express your answer numerically
Kgoal = ?

Part B
Determine the equilibrium constant, Kgoal, for the reaction
4PCl5(g)<-->P4(s)+10Cl2(g), Kgoal =?
by making use of the following information:

1. P4(s)+6Cl2(g)<-->4PCl3(g) K1 = 2.00 x10^19

2. PCl5(g)<-->PCl3(g)+Cl2(g) K2 = 1.13x10^-2

Express your answer numerically
Kgoal=?
Answer #1
A) Kgoal = (K1*K2*K3)^(0.5) = 1.714 x10^27
B) Kgoal = (K2^4)/K1 = 8.152 x10^-28